This Question: 10 pts 25 of 27 (25 complete) This Quiz: 70 pls possible hes to compare its inventory tumover to those of industry leaders, who have turnover of about 13 times per year and 8% of their assets invested in inventory Baker Míg. Inc. Net Revenue Cost of sales Inventory Total assets $27,500 $21,120 $1,260 $16,380 a) What is Bakers inventory turnover? times per year (round your response to two decimal places) b) What is Bakers percent of assets committed to i ventory? % enter your response c) How does Bakers performance compare to the industry leaders? as a percentage rounded to two decimal places) In Line with Industry Worse Better Enter your answer in each of the answer boxes. 6:47 PM
Answer #1

a) Inventory turnover = Cost of goods Sold/Inventory = 21120/1260 = 16.76%

b) Percentage of assets committed to inventory = Inventory/total assets = 1260/16380 = 7.69%

c) Better. As it has higher inventory turnover and lower assets to inventory percentage.